# The vast world of quantum advantage

### v1

@@ I The quest for quantum advantage
At first glance, Alice’s quantum approach seems unbeatable. However, Bob has a clever solution using only classical importance sampling to exactly match Alice’s quantum performance. The inner product can be estimated in classical polynomial time without any quantum technology. We also note that Alice’s ability to store the information into the amplitudes of a state in $\mathrm{poly}(n)$ is sometimes referred to as quantum random access memory (QRAM). The best known scheme for creating fault-tolerant QRAM generally requires a quantum spacetime footprint scaling as $\mathcal{O}(2^{n})$, so this advantage can disappear due to the existing challenge in physical implementations Jaques and Rattew (2023); Dalzell et al. (2025).
@@
Such pseudo-advantages have appeared repeatedly throughout the development of quantum technology. Perhaps the most dramatic story began with the HHL algorithm, a quantum algorithm for solving linear systems of equations that exemplifies genuine quantum advantage. In 2009, Harrow, Hassidim, and Lloyd proved that quantum computers can solve linear systems superpolynomially faster than classical computers Harrow et al. (2009). More precisely, they proved that this superpolynomial quantum advantage follows from the widely believed conjecture that quantum computer is strictly more powerful than classical computers in at least one computational tasks, or $\mathsf{BPP}\neq\mathsf{BQP}$. This breakthrough inspired hundreds of other derived quantum algorithms for machine learning, optimization, and numerical linear algebra, many claiming exponential speedups over classical approaches.
@@
However, starting in 2018, this excitement collided with reality. Researchers discovered that some of these derived quantum algorithms could be matched by clever classical approaches Tang (2019, 2021), similar to Bob’s solution in Puzzle I. While the HHL algorithm still has genuine quantum advantage and may provide practical value in certain applications, many previously proposed applications of the HHL algorithm for practically useful tasks turn out to offer only modest computational speedups. This pattern, apparent quantum advantages dissolving under careful analysis, teaches us a profound lesson: even when building upon genuine quantum advantages, we must rigorously verify that the advantage persists in our specific application of interest.
@@
The story of quantum advantages often takes multiple unexpected turns. Consider one of the first enticing facts one learns about in the first chapter quantum computing textbook: a quantum state of $n$ qubits contains $2^{n}$ complex amplitudes, suggesting an exponential advantage in information storage compared to classical bits. However, the next chapter usually dashes these hopes as Holevo’s bound Holevo (1973) proves that only $\mathcal{O}(n)$ classical bits of information can be reliably retrieved from $n$ qubits through measurement. Yet remarkably, this is not the end of the story. The following puzzle illustrates how quantum advantages can reemerge in subtle ways.
@@
At first glance one may conclude that quantum states offer a simple way to exponentially compress data for communication. However, Holevo’s bound appears to quickly and completely shatter this hope as Bob can only extract $\mathcal{O}(n)$ classical bits from his quantum states, suggesting Alice could have achieved the same result by sending classical bits. Yet this is not the complete story. The answer crucially depends on what Bob wants to do with the data. For some tasks, like computing the overlap with another vector, the compression into a quantum state offers no advantage Tang (2019). However, for other tasks, such as determining whether the vector belongs to a particular subspace or its orthogonal complement, compressing data into a quantum state provides a genuine exponential quantum advantage. Classical communication for solving this problem requires $\mathcal{O}(2^{n})$ bits, while quantum communication only requires $\mathcal{O}(n)$ qubits Raz (1999). Moreover, this quantum advantage does not rely on any unproven computational conjecture. This type of advantage has found recent applications in quantum machine learning, such as for computing quadratic functions of data Gilboa et al. (2024).
These examples underscore the importance of rigorous analysis of quantum advantages. As quantum technologies rapidly advance and require substantial investments of human effort and financial resources, establishing rigorous theoretical foundations becomes essential to ensure these efforts are well-directed. Quantum computing ventures now attract billions of dollars in investment, and governments worldwide launch major quantum initiatives, making the stakes for establishing genuine quantum advantages increasingly important. Without solid theoretical foundations, we risk pursuing technological directions that may ultimately prove less effective than anticipated, potentially misallocating significant resources and undermining confidence in quantum technologies. The quantum computing community continues to grapple with these challenges, debating evaluation criteria for quantum algorithms and addressing misconceptions about near-term capabilities, while recognizing the critical need for algorithmic breakthroughs to justify ongoing hardware investments Zimborás et al. (2025); Aaronson et al. (2025); King (2025); Lanes et al. (2025).
@@ II.1 Predictability
The evolution of quantum recommendation systems illustrates the subtle challenges in predicting quantum advantages. Early work Kerenidis and Prakash (2017) proposed quantum algorithms that appeared to offer exponential speedups for recommendation problems, representing a natural application area where quantum advantages seemed plausible. However, the wide array of classical algorithmic techniques made it difficult to definitively establish the advantage. Building on techniques similar to those in Puzzle I, researchers later found that classical algorithms could achieve comparable performance Tang (2019), demonstrating how some apparent quantum advantages can dissolve when the full classical algorithmic landscape is thoroughly explored.
@@
A second powerful approach proceeds via reductions to widely believed conjectures. The key is identifying conjectures supported by diverse evidence across different fields. We then prove that the absence of quantum advantage would violate these conjectures. Common examples include $\mathsf{BPP}\neq\mathsf{BQP}$ (not all quantum polynomial-time problems are classically solvable in polynomial time), the non-collapse of the polynomial hierarchy Stockmeyer (1976), and the classical hardness of factoring Shor (1997). Even quantum computational hardness assumptions, such as the hardness for quantum computers to solve the learning with errors problem Regev (2009), can sometimes be transformed into proofs of quantum advantage Mahadev (2018); Schuster et al. (2024).
Example 2 (Shor’s algorithm for factoring Shor (1997)).
Shor’s algorithm provides compelling evidence for quantum advantage through reduction to the hardness of factoring, a problem underpinning trillions of dollars of global digital commerce UN Conference on Trade and Development (2024); Forrester Research (2024). RSA encryption, standardized by NIST as a federal cryptographic standard National Institute of Standards and Technology (2013), forms a critical component of this digital infrastructure. Despite enormous financial incentives and decades of intensive research by cryptographers and computer scientists, no efficient classical algorithm has emerged. The proven ability of Shor’s algorithm to factor efficiently, combined with the sustained difficulty of classical approaches, provides strong evidence for quantum advantage.
Example 3 (Cooling physical systems Chen et al. (2024a)).
@@
Example 4 (Decoded quantum interferometry Jordan et al. (2024)).
Decoded quantum interferometry (DQI) exemplifies this third paradigm. DQI reduces classical optimization problems to classical decoding problems through quantum interference. Most remarkably, the performance of this quantum algorithm can be predicted by evaluating classical decoders using classical computers, even for large-scale instances. This enables rigorous comparison with the best classical algorithms before building quantum hardware. For certain optimization problems, such as optimal polynomial intersection, this approach has revealed that DQI can efficiently achieve better approximation ratios than any known classical algorithm. This provides evidence for quantum advantage through a combination of theoretical analysis and empirical computation.
Example 5 (Quantum phase estimation Kitaev (1995); Lin and Tong (2020); Wu et al. (2024)).
This paradigm also manifests in finding low-energy states of a Hamiltonian $H$. After classical algorithms like DMRG White (1992) find their best possible approximation to the ground state, we can predict how the quantum phase estimation algorithm Kitaev (1995); Lin and Tong (2020) would improve upon this result. This prediction uses a key quantum observable: if the energy variance $\operatorname{tr}(H^{2}\rho)-\operatorname{tr}(H\rho)^{2}$ of the classically obtained state $\rho$ is substantial, we can guarantee that quantum phase estimation will efficiently find states with lower energy using the state found classically as input Wu et al. (2024). The variance, often efficiently computable classically, provides a rigorous lower bound on how much the quantum algorithm can outperform that specific classical result. This enables precise prediction and quantification of quantum advantage in energy minimization without requiring quantum hardware.
@@ II.2 Typicality
Consider the task of simulating a closed quantum system evolving under a time-dependent Hamiltonian described by a polynomial-size quantum circuit $C$. The goal is to estimate local observables on the output state after evolving a simple product input state under circuit $C$. Under the widely believed conjecture that $\mathsf{BPP}\neq\mathsf{BQP}$, this task exhibits a superpolynomial quantum advantage in the worst case; that is, for the quantum circuits that are most difficult to simulate classically. However, this quantum advantage only applies for some quantum circuits. For most quantum circuits consisting of independent and generic circuit layers, the quantum chaos in the dynamics enables efficient classical simulation through low-weight Pauli propagation Aharonov et al. (2023); Shao et al. (2023); Nemkov et al. (2023); Begušić et al. (2024, 2023); Fontana et al. (2023); Rudolph et al. (2023); Angrisani et al. (2024). Moreover, it is widely believed that under natural Hamiltonian dynamics, local observables approach the predictions from generalized thermal equilibrium states characterized by a few conserved quantities such as particle numbers. This stark contrast between worst-case and average-case complexity exemplifies why worst-case advantages need not imply quantum advantages for typical instances.
@@
Example 7 (Random circuit sampling Arute et al. (2019); Zhu et al. (2022); Morvan et al. (2024); Abanin et al. (2025); Gao et al. (2025)).
Random circuit sampling (RCS) exemplifies how average-case hardness can establish typical quantum advantages. The task involves sampling from the output distribution of a random quantum circuit to within a specified total variation distance. For example, Google’s Willow processor demonstrated quantum supremacy by sampling from a $103$-qubit random circuit in just a few minutes, while a classical computer would require much longer than the age of the universe to complete the same task Abanin et al. (2025). The classical hardness of RCS stems from a reduction. For most circuit instances, approximate sampling to small total variation distance requires superpolynomial classical computation, under the widely believed assumption that the polynomial hierarchy does not collapse. These results establish that quantum advantages for RCS are typical. They hold for most random quantum circuits rather than carefully constructed worst-case instances.
@@
The discrete logarithm problem (DLP) provides another prototypical example of typical quantum advantage. Given a cyclic group $G$ of prime order $p$ with generator $g$, and an element $h\in G$, the task is to find $x$ such that $g^{x}=h$. While Shor’s algorithm solves DLP in quantum polynomial time, the best-known classical algorithms require exponential time for most group elements when $G$ is carefully chosen Boneh (1998). The generic nature of this quantum advantage stems from a powerful reduction: finding discrete logarithms for a random element $h\in G$ is provably as hard as solving the problem in the worst case Shoup (1997). This reduction is known as random self-reducibility and holds for broad classes of groups, including those used in practical cryptographic systems, making the average-case hardness a foundation for widely deployed cryptographic protocols.
These examples highlight how average-case complexity can establish quantum advantages that are both theoretically rigorous and practically relevant. However, developing the theory of average-case complexity remains an ongoing challenge. The field must address several fundamental questions: identifying natural probability distributions over problem instances that capture real-world applications, finding more problems with properties like random self-reducibility that connect average-case to worst-case hardness, and developing new proof techniques for average-case lower bounds on classical computation. As of now, most superpolynomial quantum advantages are not known to be typical. The most well known examples are random circuit sampling, where the randomly generated bitstrings are not known to be of much use, and Shor’s algorithm for breaking cryptographic protocols, where one can reinstate security by using post-quantum cryptography Regev (2009). More recently, the problem of optimal polynomial interpolation (OPI) has been conjectured to be another such instance. For OPI, there are parameter regimes that are efficiently solvable by DQI Jordan et al. (2024) but which are not solved in polynomial time by any known classical algorithms, even for average-case instances, though this problem is not known to have the property of random self-reducibility like DLP. Aside from such superpolynomial speedups, there are also a variety of average-case-hard problems, for example Circuit-SAT, that admit Grover-like quadratic speedups for typical instances. More recently, quantum algorithms have been discovered giving a super-quadratic polynomial speedup for the problem of sparse learning parity with noise (LPN) Schmidhuber et al. (2025). As LPN is a canonical average-case-hard problem where even a quantum computer is not expected to offer an efficient solution in general, better-than-quadratic polynomial speedups for typical instances of this problem are conceptually significant. These observations raise the following open question.
@@ II.3 Robustness
The threshold theorem ensures that a quantum circuit with $C$ perfect gates can be simulated by a noisy quantum circuit with error at most $\epsilon$, provided the noise per gate is below a threshold $p_{\text{th}}$. Traditional approaches require $\mathcal{O}(C\log^{c}(C/\epsilon))$ noisy gates for this simulation, where $c$ is a constant. Recent breakthroughs have dramatically improved this overhead: using constant-rate quantum locally testable codes, a depth-$D$ quantum circuit on $n$ qubits with $C$ ideal gates can be simulated by a noisy quantum circuit that maintains a constant space overhead (using only $\mathcal{O}(n)$ qubits) and near-logarithmic time overhead (using only $D\log^{1+o(1)}(C/\epsilon)$ depth) Gottesman (2013); Nguyen and Pattison (2024).
@@
By exploiting quantum entanglement, quantum sensors can achieve precision scaling as $1/(NT)$ using an $N$-particle probe with total sensing time $T$; this so-called Heisenberg-limited quantum enhanced scaling should be compared to the standard $1/(\sqrt{N}T)$ scaling achieved by an unentangled probe. However, this advantage dissolves in the presence of generic noise. Recent no-go theorems Zhou et al. (2018) prove that, for many sensing tasks, even small amounts of noise can completely destroy the asymptotic quantum advantage, forcing the precision to the scaling $1/\sqrt{NT}$ for both entangled and unentangled sensors. The $\sqrt{N}$ scaling advantage achieved using entangled probes in the ideal case is fundamentally unachievable when generic noise is present, even if quantum error correction is strategically deployed. A simplified, self-contained proof is given in Appendix B.
@@
Beyond hardware noise, quantum advantages must also withstand deviations from other theoretical assumptions. The problem instances occurring in the real world may not satisfy properties that guarantee quantum advantages. In physical simulation, the initial condition or the specification of the dynamics may be slightly different from those considered theoretically. And in machine learning, the classical datasets may be inherently noisy or incomplete.
In recent years, theorists have envisioned quantum artificial intelligence (AI) agents that can learn about our universe by retrieving information using quantum sensors, storing the information in quantum memory, and processing the information using quantum computers. Ideal noiseless quantum AI agents have been proven to achieve some learning tasks with exponentially fewer experiments than would be required by classical AI agents that store data in classical memory and process it with classical computers Huang et al. (2021a); Chen et al. (2022); Huang et al. (2022a). Because quantum AI agents integrate multiple quantum technologies, they face robustness challenges across multiple fronts simultaneously. The vulnerability of quantum sensing to noise may hamper data collection, while memory decoherence can corrupt stored information, and computational errors can compromise data processing. While some quantum learning advantages have been proven to persist under realistic noise conditions Huang et al. (2022a, b); Chen et al. (2023) and experimentally demonstrated in intermediate-scale superconducting qubit platforms Huang et al. (2022) and photonic systems Liu et al. (2025a), the robustness of many other learning advantages remains uncharacterized. In particular, it is not known whether universal fault-tolerant schemes exist that can preserve all exponential advantages enjoyed by ideal quantum AI agents.

### v2

@@ I The quest for quantum advantage
At first glance, Alice’s quantum approach seems unbeatable. However, Bob has a clever solution using only classical importance sampling to exactly match Alice’s quantum performance. The inner product can be estimated in classical polynomial time without any quantum technology. We also note that Alice’s ability to store the information into the amplitudes of a state in $\mathrm{poly}(n)$ is sometimes referred to as quantum random access memory (QRAM). The best known scheme for creating fault-tolerant QRAM generally requires a quantum spacetime footprint scaling as $\mathcal{O}(2^{n})$, so this advantage can disappear due to the existing challenge in physical implementations [2, 3].
@@
Such pseudo-advantages have appeared repeatedly throughout the development of quantum technology. Perhaps the most dramatic story began with the HHL algorithm, a quantum algorithm for solving linear systems of equations that exemplifies genuine quantum advantage. In 2009, Harrow, Hassidim, and Lloyd proved that quantum computers can solve linear systems superpolynomially faster than classical computers [4]. More precisely, they proved that this superpolynomial quantum advantage follows from the widely believed conjecture that quantum computer is strictly more powerful than classical computers in at least one computational tasks, or $\mathsf{BPP}\neq\mathsf{BQP}$. This breakthrough inspired hundreds of other derived quantum algorithms for machine learning, optimization, and numerical linear algebra, many claiming exponential speedups over classical approaches.
@@
However, starting in 2018, this excitement collided with reality. Researchers discovered that some of these derived quantum algorithms could be matched by clever classical approaches [5, 6], similar to Bob’s solution in Puzzle I. While the HHL algorithm still has genuine quantum advantage and may provide practical value in certain applications, many previously proposed applications of the HHL algorithm for practically useful tasks turn out to offer only modest computational speedups. This pattern, apparent quantum advantages dissolving under careful analysis, teaches us a profound lesson: even when building upon genuine quantum advantages, we must rigorously verify that the advantage persists in our specific application of interest.
@@
The story of quantum advantages often takes multiple unexpected turns. Consider one of the first enticing facts one learns about in the first chapter quantum computing textbook: a quantum state of $n$ qubits contains $2^{n}$ complex amplitudes, suggesting an exponential advantage in information storage compared to classical bits. However, the next chapter usually dashes these hopes as Holevo’s bound [7] proves that only $\mathcal{O}(n)$ classical bits of information can be reliably retrieved from $n$ qubits through measurement. Yet remarkably, this is not the end of the story. The following puzzle illustrates how quantum advantages can reemerge in subtle ways.
@@
At first glance one may conclude that quantum states offer a simple way to exponentially compress data for communication. However, Holevo’s bound appears to quickly and completely shatter this hope as Bob can only extract $\mathcal{O}(n)$ classical bits from his quantum states, suggesting Alice could have achieved the same result by sending classical bits. Yet this is not the complete story. The answer crucially depends on what Bob wants to do with the data. For some tasks, like computing the overlap with another vector, the compression into a quantum state offers no advantage [5]. However, for other tasks, such as determining whether the vector belongs to a particular subspace or its orthogonal complement, compressing data into a quantum state provides a genuine exponential quantum advantage. Classical communication for solving this problem requires $\mathcal{O}(2^{n})$ bits, while quantum communication only requires $\mathcal{O}(n)$ qubits [8]. Moreover, this quantum advantage does not rely on any unproven computational conjecture. This type of advantage has found recent applications in quantum machine learning, such as for computing quadratic functions of data [9].
These examples underscore the importance of rigorous analysis of quantum advantages. As quantum technologies rapidly advance and require substantial investments of human effort and financial resources, establishing rigorous theoretical foundations becomes essential to ensure these efforts are well-directed. Quantum computing ventures now attract billions of dollars in investment, and governments worldwide launch major quantum initiatives, making the stakes for establishing genuine quantum advantages increasingly important. Without solid theoretical foundations, we risk pursuing technological directions that may ultimately prove less effective than anticipated, potentially misallocating significant resources and undermining confidence in quantum technologies. The quantum computing community continues to grapple with these challenges, debating evaluation criteria for quantum algorithms and addressing misconceptions about near-term capabilities, while recognizing the critical need for algorithmic breakthroughs to justify ongoing hardware investments [10, 11, 12, 13].
@@ II.1 Predictability
The evolution of quantum recommendation systems illustrates the subtle challenges in predicting quantum advantages. Early work [14] proposed quantum algorithms that appeared to offer exponential speedups for recommendation problems, representing a natural application area where quantum advantages seemed plausible. However, the wide array of classical algorithmic techniques made it difficult to definitively establish the advantage. Building on techniques similar to those in Puzzle I, researchers later found that classical algorithms could achieve comparable performance [5], demonstrating how some apparent quantum advantages can dissolve when the full classical algorithmic landscape is thoroughly explored.
@@
A second powerful approach proceeds via reductions to widely believed conjectures. The key is identifying conjectures supported by diverse evidence across different fields. We then prove that the absence of quantum advantage would violate these conjectures. Common examples include $\mathsf{BPP}\neq\mathsf{BQP}$ (not all quantum polynomial-time problems are classically solvable in polynomial time), the non-collapse of the polynomial hierarchy [15], and the classical hardness of factoring [16]. Even quantum computational hardness assumptions, such as the hardness for quantum computers to solve the learning with errors problem [17], can sometimes be transformed into proofs of quantum advantage [18, 19].
Example 2 (Shor’s algorithm for factoring [16]).
Shor’s algorithm provides compelling evidence for quantum advantage through reduction to the hardness of factoring, a problem underpinning trillions of dollars of global digital commerce [20, 21]. RSA encryption, standardized by NIST as a federal cryptographic standard [22], forms a critical component of this digital infrastructure. Despite enormous financial incentives and decades of intensive research by cryptographers and computer scientists, no efficient classical algorithm has emerged. The proven ability of Shor’s algorithm to factor efficiently, combined with the sustained difficulty of classical approaches, provides strong evidence for quantum advantage.
Example 3 (Cooling physical systems [23]).
@@
Example 4 (Quantum approximate optimization algorithm [24]).
Recent advances in understanding the Quantum Approximate Optimization Algorithm (QAOA) [24] illustrate this paradigm. For specific problem classes, QAOA’s performance can be predicted through classical numerical computation [25, 26, 27]. Remarkably, rigorous lower bounds can be established even when running the quantum algorithm is infeasible with current technology. For MaxCut on high-girth 3-regular graphs, this analysis has provided evidence for potential quantum advantages over known classical algorithms [27].
Example 5 (Decoded quantum interferometry [28]).
Decoded quantum interferometry (DQI) [28] provides another compelling illustration. DQI reduces classical optimization problems to classical decoding problems through quantum interference. DQI’s performance can be predicted by evaluating classical decoders, even for large-scale instances. This enables comparison with classical algorithms before building quantum hardware. For certain problems, such as optimal polynomial intersection, this approach has revealed that DQI achieves better approximation ratios than any known classical algorithm [28].
Example 6 (Quantum phase estimation [29, 30, 31]).
This paradigm also manifests in quantum phase estimation (QPE) for finding low-energy states of a Hamiltonian $H$. After classical algorithms like DMRG [32] find their best approximation to the ground state, we can predict how quantum phase estimation would improve upon this result. If the energy variance $\operatorname{tr}(H^{2}\rho)-\operatorname{tr}(H\rho)^{2}$ of the classical state $\rho$ is substantial, QPE can efficiently find lower-energy states [31]. This variance, often efficiently computable classically, provides a rigorous lower bound on the quantum advantage without requiring quantum hardware.
@@ II.2 Typicality
Consider the task of simulating a closed quantum system evolving under a time-dependent Hamiltonian described by a polynomial-size quantum circuit $C$. The goal is to estimate local observables on the output state after evolving a simple product input state under circuit $C$. Under the widely believed conjecture that $\mathsf{BPP}\neq\mathsf{BQP}$, this task exhibits a superpolynomial quantum advantage in the worst case; that is, for the quantum circuits that are most difficult to simulate classically. However, this quantum advantage only applies for some quantum circuits. For most quantum circuits consisting of independent and generic circuit layers, the quantum chaos in the dynamics enables efficient classical simulation through low-weight Pauli propagation [33, 34, 35, 36, 37, 38, 39, 40, 41]. Moreover, it is widely believed that under natural Hamiltonian dynamics, local observables approach the predictions from generalized thermal equilibrium states characterized by a few conserved quantities such as particle numbers. This stark contrast between worst-case and average-case complexity exemplifies why worst-case advantages need not imply quantum advantages for typical instances.
@@
Example 8 (Random circuit sampling [42, 43, 44, 45, 46]).
Random circuit sampling (RCS) exemplifies how average-case hardness can establish typical quantum advantages. The task involves sampling from the output distribution of a random quantum circuit to within a specified total variation distance. For example, Google’s Willow processor demonstrated quantum supremacy by sampling from a $103$-qubit random circuit in just a few minutes, while a classical computer would require much longer than the age of the universe to complete the same task [45]. The classical hardness of RCS stems from a reduction. For most circuit instances, approximate sampling to small total variation distance requires superpolynomial classical computation, under the widely believed assumption that the polynomial hierarchy does not collapse. These results establish that quantum advantages for RCS are typical. They hold for most random quantum circuits rather than carefully constructed worst-case instances.
@@
The discrete logarithm problem (DLP) provides another prototypical example of typical quantum advantage. Given a cyclic group $G$ of prime order $p$ with generator $g$, and an element $h\in G$, the task is to find $x$ such that $g^{x}=h$. While Shor’s algorithm solves DLP in quantum polynomial time, the best-known classical algorithms require exponential time for most group elements when $G$ is carefully chosen [47]. The generic nature of this quantum advantage stems from a powerful reduction: finding discrete logarithms for a random element $h\in G$ is provably as hard as solving the problem in the worst case [48]. This reduction is known as random self-reducibility and holds for broad classes of groups, including those used in practical cryptographic systems, making the average-case hardness a foundation for widely deployed cryptographic protocols.
These examples highlight how average-case complexity can establish quantum advantages that are both theoretically rigorous and practically relevant. However, developing the theory of average-case complexity remains an ongoing challenge. The field must address several fundamental questions: identifying natural probability distributions over problem instances that capture real-world applications, finding more problems with properties like random self-reducibility that connect average-case to worst-case hardness, and developing new proof techniques for average-case lower bounds on classical computation. As of now, most superpolynomial quantum advantages are not known to be typical. The most well known examples are random circuit sampling, where the randomly generated bitstrings are not known to be of much use, and Shor’s algorithm for breaking cryptographic protocols, where one can reinstate security by using post-quantum cryptography [17]. More recently, the problem of optimal polynomial interpolation (OPI) has been conjectured to be another such instance. For OPI, there are parameter regimes that are efficiently solvable by DQI [28] but which are not solved in polynomial time by any known classical algorithms, even for average-case instances, though this problem is not known to have the property of random self-reducibility like DLP. Aside from such superpolynomial speedups, there are also a variety of average-case-hard problems, for example Circuit-SAT, that admit Grover-like quadratic speedups for typical instances. More recently, quantum algorithms have been discovered giving a super-quadratic polynomial speedup for the problem of sparse learning parity with noise (LPN) [49]. As LPN is a canonical average-case-hard problem where even a quantum computer is not expected to offer an efficient solution in general, better-than-quadratic polynomial speedups for typical instances of this problem are conceptually significant. These observations raise the following open question.
@@ II.3 Robustness
The threshold theorem ensures that a quantum circuit with $C$ perfect gates can be simulated by a noisy quantum circuit with error at most $\epsilon$, provided the noise per gate is below a threshold $p_{\text{th}}$. Traditional approaches require $\mathcal{O}(C\log^{c}(C/\epsilon))$ noisy gates for this simulation, where $c$ is a constant. Recent breakthroughs have dramatically improved this overhead: using constant-rate quantum locally testable codes, a depth-$D$ quantum circuit on $n$ qubits with $C$ ideal gates can be simulated by a noisy quantum circuit that maintains a constant space overhead (using only $\mathcal{O}(n)$ qubits) and near-logarithmic time overhead (using only $D\log^{1+o(1)}(C/\epsilon)$ depth) [50, 51].
@@
By exploiting quantum entanglement, quantum sensors can achieve precision scaling as $1/(NT)$ using an $N$-particle probe with total sensing time $T$; this so-called Heisenberg-limited quantum enhanced scaling should be compared to the standard $1/(\sqrt{N}T)$ scaling achieved by an unentangled probe. However, this advantage dissolves in the presence of generic noise. Recent no-go theorems [52] prove that, for many sensing tasks, even small amounts of noise can completely destroy the asymptotic quantum advantage, forcing the precision to the scaling $1/\sqrt{NT}$ for both entangled and unentangled sensors. The $\sqrt{N}$ scaling advantage achieved using entangled probes in the ideal case is fundamentally unachievable when generic noise is present, even if quantum error correction is strategically deployed. A simplified, self-contained proof is given in Appendix B.
@@
Beyond hardware noise, quantum advantages must also withstand deviations from other theoretical assumptions. The problem instances occurring in the real world may not satisfy properties that guarantee quantum advantages. In physical simulation, the initial condition or the specification of the dynamics may be slightly different from those considered theoretically. And in machine learning, the classical datasets may be inherently noisy or incomplete [53, 54, 55].
In recent years, theorists have envisioned quantum artificial intelligence (AI) agents that can learn about our universe by retrieving information using quantum sensors, storing the information in quantum memory, and processing the information using quantum computers. Ideal noiseless quantum AI agents have been proven to achieve some learning tasks with exponentially fewer experiments than would be required by classical AI agents that store data in classical memory and process it with classical computers [56, 57, 58]. Because quantum AI agents integrate multiple quantum technologies, they face robustness challenges across multiple fronts simultaneously. The vulnerability of quantum sensing to noise may hamper data collection, while memory decoherence can corrupt stored information, and computational errors can compromise data processing. While some quantum learning advantages have been proven to persist under realistic noise conditions [58, 59, 60] and experimentally demonstrated in intermediate-scale superconducting qubit platforms [58] and photonic systems [61], the robustness of many other learning advantages remains uncharacterized. In particular, it is not known whether universal fault-tolerant schemes exist that can preserve all exponential advantages enjoyed by ideal quantum AI agents.